Find the quotient.
y*9 ÷ y*3
hear are the answers
A)y1/3
B)y*3
C)y*6
D)y*12

Answer by stanbon(75887)

Find the quotient.
y*9 ÷ y*3 = (yyyyyyyyy)/(yyy) = yyyyyy = y^(9-3) = y^6
==================================
Cheers,
Stan H.
==================================
hear are the answers
A)y1/3
B)y*3
C)y*6
D)y*12